# Kai Zeng    Zhanqian Wu    Kaixin Xiong    Xiaobao Wei    Xiangyu Guo    Zhenxin Zhu    Kalok Ho    Lijun Zhou    Bohan Zeng    Ming Lu    Haiyang Sun    Bing Wang    Guang Chen    Hangjun Ye    Wentao Zhang [ [ [

### v3

@@ 5.2 Main Results
1x Epochs 2x Epochs
(28130) (+420, <2%) (28130) (wang2024drivedreamer) (lu2024wovogen) (gao2023magicdrive) (wen2024panacea) (huang2024subjectdrive) (Ours)
mAP $\uparrow$ 34.5 36.1 38.4 35.8 36.2 35.4 37.1 38.1 38.7
mAVE $\downarrow$ 29.1 28.9 27.7 123.4 27.3 26.4 26.8
NDS $\uparrow$ 46.9 47.8 50.4 39.5 18.1 39.8 49.2 50.2 50.6
Table 1: Comparison of detection under different training epochs. * indicates the evaluation of WoVoGen is only on the vehicle classes of cars, trucks, and buses. Bold and underlines indicate the best and second best results respectively.
1x Epochs 2x Epochs
Real Dream4Drive Real Panacea SubjectDrive Dream4Drive
(28130) (+420, <2%) (28130) (wen2024panacea) (huang2024subjectdrive) (Ours)
AMOTA $\uparrow$ 30.1 31.2 34.1 33.7 33.7 34.4
AMOTP $\downarrow$ 137.9 135.4 134.1 135.3 135.3 133.5
Table 2: Comparison of tracking under different training epochs. Bold and underlines indicate the best and second best.
@@
As generative models continue to advance, the ability to synthesize high-resolution videos has become achievable (gao2024vista; gao2024magicdrivedit). To investigate the effect of high-resolution synthetic data on downstream perception models, we further conduct experiments for detection and tracking tasks at a resolution of $512\times 768$, as reported in Tab. 3 and 4.
@@
Table 3: Detection performance under different training epochs (1x, 2x, 3x). “Naive Insert" denotes the direct projection of 3D assets into the original scene. Results are reported at 512×768 resolution.
Quantitative and Qualitative Comparison with Naive Insertion.
@@
Table 4: Tracking performance under different training epochs (1x, 2x, 3x). “Naive Insert" denotes the direct projection of 3D assets into the original scene. Results are reported at 512×768 resolution.
Comprehensive visualization of asset insertion results.
@@ 5.3 Ablation Studies
We observe that the source of inserted assets affects the quality of synthetic data, consistent with (ljungbergh2025r3d2). We therefore investigate how different asset sources influence downstream perception performance.
Views Distances 3D Asset Generation Methods
Front Back Left Right Close Mid Far Trellis Hunyuan3D Ours
mAP $\uparrow$ 40.2 40.2 40.2 39.8 39.7 40.3 40.5 39.8 40.2 40.7
mATE $\downarrow$ 66.2 66.0 64.6 66.2 65.7 65.4 65.1 65.6 65.1 64.2
mAOE $\downarrow$ 51.2 55.1 45.7 51.4 52.2 51.7 49.7 51.8 50.8 48.0
mAVE $\downarrow$ 27.7 27.5 28.5 27.9 28.1 28.0 27.9 27.8 28.0 27.1
NDS $\uparrow$ 51.0 50.6 51.6 50.7 50.5 50.9 51.3 50.8 50.9 52.0
@@
As shown in Tab. 5, although Trellis (xiang2025structured) can generate high-quality assets, its style does not fully match autonomous driving scenarios, which can lead to artifacts and degraded quality when assets are inserted, negatively affecting downstream tasks. Hunyuan3D (hunyuan3d2025hunyuan3d)’s single-view generation also underperforms our multiview approach, since single-image assets may be incomplete, whereas our method produces complete, high-quality 3D assets.
@@ 6 Conclusion
Appendix
@@ A.1 Model and Training Details
DriveObj3D Generation Pipeline. We adopt Grounded-SAMren2024grounded for image segmentation, Qwen-Image-Editwu2025qwenimagetechnicalreport for image generation, and Hunyuan3D 2.0hunyuan3d2025hunyuan3d for 3D asset generation.
Inpainting Diffusion Model. For video synthesis, Dream4Drive is constructed upon a DiT-based architecture. The backbone is initialized with pretrained weights from MagicDriveDiT (gao2024magicdrivedit), which originally conditioned on BEV maps and 3D bounding boxes. We extend its ControlNet branch by incorporating novel 3D-aware guidance maps, and fine-tune the model on the nuScenes dataset. The generated videos have a resolution of $512\times 768$ with 33 frames.
@@
LPIPS Loss. To further improve perceptual quality, we adopt the LPIPS (Learned Perceptual Image Patch Similarity) loss (zhang2018unreasonable), which measures feature-level perceptual differences between prediction and ground truth:
@@
A.2 Experiment Details
Scene Insertion. We select video frames where no other vehicles appear along the insertion trajectory. For training downstream perception models, inserted scenes are strictly drawn from the nuScenes training split.
Asset Insertion. In the setting of Sec. 5.2, we use 420 samples evenly distributed across object categories, which provide the necessary coverage while keeping the synthetic data below 2% of the training set.
Ablation Studies. Experiments on insertion direction, insertion distance, and asset source in Sec. 5.3 are conducted under the 1× training epoch and 512×768 resolution setting.
@@ Appendix B Detailed AP Metrics
Tab. 3 compares detection metrics across different training epochs. Detailed per-class AP scores are provided in Tab. 6, 7, and 8.
@@ Appendix C Additional Experiments on Generation Quality
DriveDream-2 zhao2025drivedreamer 12Hz $\times512$ 55.70 11.20
@@
[…]
No cutout&mask 66.36 14.38
No depth&normal 34.64 7.33
No edge 49.45 8.44
Table 11: Ablation study on 3D-aware guidance maps.
Controllability. The controllability of our approach is quantitatively evaluated using perception metrics from StreamPETR wang2023exploring. We first generate the entire nuScenes validation set with Dream4Drive, and then measure perception performance using a pre-trained StreamPETR model. The relative metrics, compared to those obtained on real data, indicate how well the generated samples align with the original annotations. As shown in Tab. 9, Dream4Drive achieves a relative performance of 82%, demonstrating precise control over foreground object locations.
Generation Quality. As shown in Tab. 10, our method achieves FVD 31.84 and FID 5.80, outperforming layout- and 3D semantics-guided methods gao2023magicdrive; gao2024magicdrivedit; li2024uniscene; ji2025cogen. The results indicate improved motion consistency and preserved visual fidelity, confirming that our 3D-aware guidance maps effectively generate both foreground and background content. Notably, no post-processing or selective filtering was applied to the generated videos.
Appendix D Additional Visualization
Appendix E Limitation and Future Works
Appendix F Statement on LLM Usage

### v4

@@ 5.2 Main Results
As generative models continue to advance, the ability to synthesize high-resolution videos has become achievable (Gao et al., 2024b; a). To investigate the effect of high-resolution synthetic data on downstream perception models, we further conduct experiments for detection and tracking tasks at a resolution of $512\times 768$, as reported in Tab. 3 and 4.
Quantitative and Qualitative Comparison with Naive Insertion.
Comprehensive visualization of asset insertion results.
@@
Table 3: Detection performance under different training epochs (1x, 2x, 3x). “Naive Insert” denotes the direct projection of 3D assets into the original scene. Results are reported at 512×768 resolution.
@@
Table 4: Tracking performance under different training epochs (1x, 2x, 3x). “Naive Insert” denotes the direct projection of 3D assets into the original scene. Results are reported at 512×768 resolution.
@@ 5.3 Ablation Studies
We observe that the source of inserted assets affects the quality of synthetic data, consistent with (Ljungbergh et al., 2025). We therefore investigate how different asset sources influence downstream perception performance.
As shown in Tab. 5, although Trellis (Xiang et al., 2025) can generate high-quality assets, its style does not fully match autonomous driving scenarios, which can lead to artifacts and degraded quality when assets are inserted, negatively affecting downstream tasks. Hunyuan3D (Hunyuan3D et al., 2025)’s single-view generation also underperforms our multiview approach, since single-image assets may be incomplete, whereas our method produces complete, high-quality 3D assets.
Views Distances 3D Asset Generation Methods Speed
Front Back Left Right Close Mid Far Trellis Hunyuan3D Ours Low Middle High
mAP $\uparrow$ 40.2 40.2 40.2 39.8 39.7 40.3 40.5 39.8 40.2 40.7 40.4 40.7 40.9
mATE $\downarrow$ 66.2 66.0 64.6 66.2 65.7 65.4 65.1 65.6 65.1 64.2 64.5 64.2 63.9
mAOE $\downarrow$ 51.2 55.1 45.7 51.4 52.2 51.7 49.7 51.8 50.8 48.0 48.3 48.1 48.0
mAVE $\downarrow$ 27.7 27.5 28.5 27.9 28.1 28.0 27.9 27.8 28.0 27.1 27.1 27.1 27.0
NDS $\uparrow$ 51.0 50.6 51.6 50.7 50.5 50.9 51.3 50.8 50.9 52.0 51.8 52.0 52.1
@@ 6 Conclusion
Acknowledgments
This work is supported by National Natural Science Foundation of China (92470121, 62402016), National Key R&D Program of China (2024YFA1014003), Zhongguancun Academy, (Grant No.s C20250204, C20250602), and High-performance Computing Platform of Peking University.
Ethics Statement
This work focuses on synthetic data generation for autonomous driving, aiming to improve downstream perception through the proposed Dream4Drive framework. We build upon publicly available driving datasets and a curated 3D asset library, without collecting new human-subject data or personally identifiable information. Potential risks include the misuse of synthetic driving scenarios for surveillance or malicious applications, as well as biases introduced by the underlying datasets and assets. We encourage responsible use of this research strictly for advancing safe and reliable autonomous driving.
Reproducibility Statement
We ensure reproducibility by providing detailed descriptions of algorithms, metrics, and experimental settings in the main paper and appendix. All assets and code will be released upon publication.
@@ A.1 Model and Training Details
DriveObj3D Generation Pipeline. We adopt Grounded-SAM for image segmentation, Qwen-Image-Edit for image generation, and Hunyuan3D 2.0 for 3D asset generation.
Inpainting Diffusion Model. For video synthesis, Dream4Drive is constructed upon a DiT-based architecture. The backbone is initialized with pretrained weights from MagicDriveDiT (Gao et al., 2024a), which originally conditioned on BEV maps and 3D bounding boxes. We extend its ControlNet branch by incorporating novel 3D-aware guidance maps, and fine-tune the model on the nuScenes dataset. The generated videos have a resolution of $512\times 768$ with 33 frames.
@@
LPIPS Loss. To further improve perceptual quality, we adopt the LPIPS (Learned Perceptual Image Patch Similarity) loss (Zhang et al., 2018), which measures feature-level perceptual differences between prediction and ground truth:
@@
A.2 Detailed synthesis procedure and time-cost analysis for generating the 420 samples
We begin by clarifying the construction of the 420 samples, then present a detailed, step-by-step account of the time required, and conclude by describing the operations carried out at each stage.
Construction of the 420 samples. We randomly select seven annotated scenes and insert ten object categories (one asset per category). Each video contains 33 frames, from which we extract six key frames for downstream training, resulting in a total of $7\times 10\times 6=420$ samples.
Step Automation (s) Manual effort (s)
Scene selection & initial position annotation (MLLM-assisted) $700\times 10$
Scene verification $7\times 20$
Scene orientation annotation $7\times 15$
Asset selection $10\times 5$
Video generation (DiT) $70\times 10\times 4$
Total $9800\text{s}\approx 2.72\text{h}$ $295\text{s}$
Table 6: Automation and manual efforts involved in synthesizing the 420 samples.
Detailed synthesis steps and time analysis (Tab. 6.)
  • Scene selection and initial position annotation. An MLLM identifies scenes containing at least one collision-free straight-driving trajectory (front, rear, left, or right). For eligible scenes, it annotates the asset’s initial insertion position in ego-vehicle coordinates $(x,y,z)$ for the first frame. The nuScenes training set contains 700 videos, and each MLLM query takes about 10 seconds, resulting in a total of $700\times 10$ seconds. The full prompt is provided in Listing LABEL:lst:driving_prompt
  • Human verification of MLLM-filtered scenes. Only seven scenes are needed. We manually verify the MLLM-generated initial positions, requiring approximately 20 seconds per scene, totaling $7\times 20$ seconds.
  • Scene orientation annotation. Orientation is calibrated using the pyrender package. All assets are pre-standardized to ensure consistent orientation. For each scene, we render an initial asset insertion using nuScenes intrinsics and extrinsics to determine the correct yaw angle. The first rendering takes 6 seconds, the correction takes 3 seconds, and the second rendering takes another 6 seconds, totaling roughly 15 seconds per scene (i.e., $7\times 15$ seconds).
  • Asset selection. We require only one asset per object category. A random choice per category suffices (about 5 seconds each), giving a total of $10\times 5$ seconds.
  • Video generation. We generate 70 videos for the 420 samples. Each video uses 10 inference steps; on an H200 GPU each step takes roughly 4 seconds, producing a total cost of $70\times 10\times 4$ seconds.
Summary.
We adopt a partially automated pipeline. The full synthesis process for all 420 samples takes under 3 hours, with approximately 300 seconds of actual human labor, which is effectively negligible.
A.3 More Experiment Results
Quantitative Evaluation of Asset Quality. We evaluate synthesized 3D assets by measuring image similarity to their corresponding ground-truth originals using CLIP and DINO metrics (Fig. 7). For CLIP model, we adopt clip-vit-large-patch14-336 Radford et al. (2021), and for DINO model, we use dinov2-base Oquab et al. (2023). The results in Tab. 7 show that our pipeline consistently outperforms Trellis and Hunyuan3D across all object categories.
car truck bus trailer construction vehicle pedestrian motorcycle bicycle traffic cone barrier
Trellis 0.8913 0.9160 0.8930 0.9041 0.8870 0.8753 0.9045 0.9273 0.9254 0.9270
CLIP IS Hunyuan3D 0.8929 0.9043 0.9031 0.9022 0.8787 0.8863 0.9121 0.9170 0.9288 0.9101
Ours 0.9048 0.9291 0.9197 0.9270 0.8996 0.9061 0.9201 0.9351 0.9305 0.9310
Trellis 0.7008 0.7266 0.7012 0.6974 0.6852 0.6721 0.6990 0.7053 0.7275 0.7304
DINO IS Hunyuan3D 0.7360 0.7631 0.7365 0.7344 0.7204 0.7074 0.7369 0.7405 0.7610 0.7652
Ours 0.8117 0.7979 0.8082 0.8040 0.7960 0.7835 0.8065 0.8129 0.8365 0.8394
Table 7: Quantitative evaluation of synthesized asset quality using CLIP Image Similarity (CLIP IS) and DINO Image Similarity (DINO IS). Higher scores indicate closer resemblance to the ground-truth assets.
These results align with the downstream perception improvements demonstrated earlier in Tab. 5, confirming the value of high-fidelity asset synthesis.
Scaling Behavior of OOD Samples. We further study how downstream performance changes when increasing the amount of OOD synthetic data. As shown in Tab. 8, expanding OOD samples from 7 to 35 scenes does not yield proportional improvements. In fact, too many OOD samples slightly degrade performance, likely because excessive OOD content dilutes in-distribution signal. Small amounts of OOD data improve robustness (Tab. 3), but aggressive scaling is not beneficial.
[…]
7 Scenes 14 Scenes 35 Scenes 7 Scenes 14 Scenes 35 Scenes 7 Scenes 14 Scenes 35 Scenes
mAP $\uparrow$ 40.7 40.4 39.7 43.6 43.1 42.3 44.5 44.1 43.6
mATE $\downarrow$ 64.2 64.4 64.5 61.6 62.0 62.5 59.8 59.5 60.3
mAOE $\downarrow$ 48.0 49.7 51.6 39.4 39.8 40.2 40.1 40.3 40.4
mAVE $\downarrow$ 27.1 27.8 28.4 27.4 28.1 28.4 27.2 27.7 27.6
NDS $\uparrow$ 52.0 51.6 50.9 54.3 53.8 53.1 55.0 54.6 54.2
Table 8: Scaling analysis with increasing numbers of OOD scenes under different training epochs. More OOD samples do not necessarily yield better performance.
To examine whether more diverse OOD conditions (rather than more scenes) provide additional gains, we apply style transfer to generate new synthetic data (e.g., rain or nighttime) and augment the original set. Tab. 9 shows consistent improvement across all metrics.
[…]
Original (420) Style Transfer (420+420) Original (420) Style Transfer (420+420) Original (420) Style Transfer (420+420)
mAP $\uparrow$ 40.7 41.2 43.6 44.2 44.5 44.8
mATE $\downarrow$ 64.2 63.7 61.6 61.2 59.8 59.7
mAOE $\downarrow$ 48.0 47.5 39.4 39.0 40.1 39.8
mAVE $\downarrow$ 27.1 26.8 27.4 26.8 27.2 26.9
NDS $\uparrow$ 52.0 52.4 54.3 54.7 55.0 55.2
Table 9: Effect of style-transferred OOD synthetic data. Adding environmental diversity improves all downstream perception metrics.
Conclusion. Our findings indicate that downstream perception benefits most from OOD scene layouts and from environmental diversity, rather than merely increasing the volume of synthetic data. Additional visualization examples of style transfer are included in the Fig. 20,  19,  21,  22.
A.4 More Ablation Studies
[…]
Ablation on Trajectory Speed. Inserted vehicles always follow straight trajectories, and we vary speed across three levels (2 / 5 / 8 m/s). Higher speeds consistently yield larger downstream gains (Tab. 5), consistent with Table 5 in the main paper. Faster inserted vehicles leave the ego-vehicle’s vicinity earlier, helping the perception model learn to detect distant objects more effectively.
Ablation on Asset Categories. For each object category, we randomly select three assets and split them into three groups (A / B / C). Performance is nearly identical across groups (Tab. 10), demonstrating that visual style differences across assets have negligible influence as long as the OOD layout is preserved.
car truck bus trailer construction vehicle pedestrian motorcycle bicycle traffic cone barrier
Assets group A 0.600 0.354 0.341 0.106 0.135 0.468 0.402 0.411 0.626 0.565
Assets group B 0.594 0.350 0.343 0.110 0.135 0.471 0.405 0.414 0.625 0.563
Assets group C 0.601 0.352 0.341 0.108 0.133 0.470 0.401 0.411 0.626 0.565
Table 10: Ablation on selected asset categories, reporting AP metrics.
Ablation on Scene Selection Diversity. We randomly sample two additional sets of insertion scenes, forming three groups (A / B / C) including the original. Downstream results remain similar across groups (Tab. 11), indicating that scene-selection-induced bias is minimal.
[…]
A B C A B C A B C
mAP $\uparrow$ 40.7 40.8 40.7 43.6 43.5 43.6 44.5 44.3 44.5
mATE $\downarrow$ 64.2 64.1 64.2 61.6 61.5 61.4 59.8 59.8 59.9
mAOE $\downarrow$ 48.0 48.0 48.1 39.4 39.4 39.5 40.1 40.2 40.1
mAVE $\downarrow$ 27.1 26.9 27.1 27.4 27.5 27.3 27.2 27.3 27.2
NDS $\uparrow$ 52.0 52.1 52.1 54.3 54.2 54.3 55.0 55.0 55.1
Table 11: Ablation on insertion-scene selection.
@@ Appendix B Detailed AP Metrics
Tab. 3 compares detection metrics across different training epochs. Detailed per-class AP scores are provided in Tab. 12, 13, and 14.
@@ Appendix C Additional Experiments on Generation Quality
[…]
No cutout&mask 66.36 14.38
No depth&normal 34.64 7.33
No edge 49.45 8.44
Table 15: Ablation study on 3D-aware guidance maps.
Controllability. The controllability of our approach is quantitatively evaluated using perception metrics from StreamPETR Wang et al. (2023a). We first generate the entire nuScenes validation set with Dream4Drive, and then measure perception performance using a pre-trained StreamPETR model. The relative metrics, compared to those obtained on real data, indicate how well the generated samples align with the original annotations. As shown in Tab. 16, Dream4Drive achieves a relative performance of 82%, demonstrating precise control over foreground object locations.
Generation Quality. As shown in Tab. 17, our method achieves FVD 31.84 and FID 5.80, outperforming layout- and 3D semantics-guided methods Gao et al. (2023; 2024a); Li et al. (2024a); Ji et al. (2025). The results indicate improved motion consistency and preserved visual fidelity, confirming that our 3D-aware guidance maps effectively generate both foreground and background content. Notably, no post-processing or selective filtering was applied to the generated videos.
Appendix D Additional Visualization
Appendix E Limitation and Future Works
Appendix F Statement on LLM Usage
@@
DriveDream-2 Zhao et al. (2025a) 12Hz $\times512$ 55.70 11.20
@@
You are an expert in autonomous-driving scene understanding.
Based on the input video, perform the following tasks without any bias or preference toward specific scene types (e.g., do not favor simple scenes over complex ones, and do not assume safety based on convenience). All decisions must be strictly derived from observable evidence in the video.
[Task Objective]
Evaluate the driving scene and determine whether there exists at least one safe, unobstructed, straight driving trajectory in any of the four directions relative to the ego vehicle: front, rear, left, or right.
If such a direction exists, identify the direction and provide the recommended initial insertion position for a new 3D asset. The position must be given in the ego-vehicle coordinate system and must correspond to the first frame of the video.
[Ego-Vehicle Coordinate System]
- The ego vehicle is located at the origin (0, 0, 0)
- x-axis: forward
- y-axis: left
- z-axis: upward
[Definition of "Safe Straight Trajectory"]
A direction is considered valid for 3D-asset insertion only if all the following conditions are met:
1. Throughout the entire video, that direction remains free of vehicles, pedestrians, and any static or dynamic obstacles blocking straight motion.
2. The road in that direction is continuous and drivable (not leading into oncoming traffic lanes, sidewalks, barriers, buildings, etc.).
3. There is sufficient spatial clearance for inserting a typical vehicle-sized 3D asset (approx. length 4m, width 1.8m, height 1.5m).
4. The insertion location must not collide with any existing object in the first frame.
5. No bias is allowed - judge each direction solely based on objective video evidence, without assuming simplicity or difficulty, and without preferring any type of scene.
[Output Format]
Respond strictly in the following JSON structure:
"is_insertable": true/false,
"direction": "front" | "rear" | "left" | "right" | null,
"initial_position": {
"x": number or null,
"y": number or null,
"z": number or null
"reason": "Brief explanation of the decision"
[Examples]
If the front direction is clear and drivable:
"is_insertable": true,
"direction": "front",
"initial_position": {"x": 8.0, "y": 0.0, "z": 0.0},
"reason": "The front lane is unobstructed and drivable for a long distance."
If no direction is suitable:
"is_insertable": false,
"direction": null,
"initial_position": {"x": null, "y": null, "z": null},
"reason": "All directions contain obstacles that block a safe straight trajectory."
Listing 1: Prompt for instructing MLLM to identify safe driving scenarios.
@@
Figure 19: Style transfer. Insertion of a bus in the rain.
Figure 20: Style transfer. Insertion of a truck in the rain.
Figure 21: Style transfer. Insertion of a barrier in the night.
Figure 22: Style transfer. Insertion of a constructive vehicle in the night.